what is the probability that at least five out of 20 young executives practice good reading habits

1
Expert's answer
2022-04-20T14:45:07-0400

Assume that there are two possibilities for each of "20" young executives: a person practices good reading habits or not. There are "2^{20}" different possibilities. The number "2^{20}" is obtained by using the multiplication principle of combinatorics. Denote by "N" a number of young executives that practice good reading habits. "N" is a random variable. We assume that chances to practice good reading habits or not for all persons are equal. The task is to find "P(N\geq5)"; "P(N\geq5)=1-P(N<5)=1-\sum_{i=0}^4P(N=i)."

"P(N=i)=\frac{C_{20}^i}{2^{20}};" "C_{20}^i=\frac{20!}{i!(20-i)!};" "C_{20}^i" is a binomial coefficient.

We get: "P(N\geq5)=\frac{2^{20}-C_{20}^0-C_{20}^1-C_{20}^2-C_{20}^3-C_{20}^4}{2^{20}}\approx0.994".

Thus, the probability that at least five persons practice good reading habits is: "0.994."